Operations Report – April 29th

Crew 193 Operations Report 29Apr2018

SOL: 7

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).
I will repair charging plugs once we have a soldering iron.
One piece suit #10’s fans stopped when I bent over while wearing the suit. The loose connection was determined to be at the positive terminal of the battery; upon inspection, there was a cold solder joint on a terminal connector. While removing tape to access battery terminals, I broke the small wires on negative terminal. Suit #10 is out of service until we have a working soldering iron (should be on the way).

Generator (hours run): 10:45 pm to 7:35 am.

Solar— SOC 57 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 2.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 455 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134315.1

Toilet tank emptied: yes
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.3

ATV Oil Checked?: Yes

ATV Oil Added?: No

ATV Fuel Used: 2.5 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda oil is within range, but on low end.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: Sawhorses moved in from RAM.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Would like some new crimp-on connectors for suit battery terminals. For example: https://www.mcmaster.com/#7060K38. An assortment would probably be nice for future missions.

Copyright © The Mars Society. All rights reserved. | Main Site